# Addressing Model and Data Heterogeneity in Multimodal Large Language Model Training

### v1

@@ 7.2.2 Data-aware Reordering
MMScale’s data-aware reordering algorithm executes before each training iteration. Table 3 presents a detailed analysis of the runtime overhead of the reordering algorithm. The DP size is the number of data parallelism groups in modality encoder. It completes the reordering in just a few milliseconds, a duration that is negligible compared to the several seconds to tens of seconds required for one iteration.
@@ 7.3 Case Study
In this subsection, we first evaluate MMScale under different frozen training settings. We then evaluate the effectiveness of system optimizations tailored for multimodal LLM training.
@@ 7.3.1 Frozen Training
We conduct a frozen training experiment under four specific training settings: complete module freezing, exclusive encoder training, exclusive LLM training, and exclusive generator training. We keep training the projectors. In these scenarios, frozen modules neither compute weight gradients nor update weights. All other experimental setup aligns with those detailed in §7.2. We evaluate the MFU and training throughput (TPT.) of MMScale compared to Megatron-LM. The experimental results are presented in Figure 19 and Figure 20. MMScale consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4%-2.9% higher MFU and 1.2%-2.9% increased training throughput. This pronounced performance gap underscores the challenges posed by Megatron-LM’s rigid resource allocation in complex training environments. In contrast, MMScale adaptively adjusts resource allocation based on training settings and consistently achieves high resource utilization.
@@
| Model | DP | Global | Intra-microbatch | Inter-microbatch |
| Size | Batch Size | Reordering | Reordering | |
| MLLM-72B | 30 | 1920 | 1.82ms | 5.87ms |
| MLLM-72B | 60 | 1920 | 1.87ms | 3.02ms |
| MLLM-72B | 120 | 1920 | 1.89ms | 1.71ms |
@@
7.3.2 System Optimizations
Asynchronous Data Loader. We conduct an experiment to evaluate the effectiveness of asynchronous data loading optimization. Setting the DP size to one, we measure the average data loading time per iteration solely on the CPU. We then compare data loading times with and without this optimization and use varying numbers of images and different image resolutions for one training iteration. The results, depicted in Figure 21, indicate that asynchronous data loading significantly reduces loading time from seconds to milliseconds. The first parameter in the x-axis represents the number of images, while the second parameter denotes the image resolution. This reduction renders the loading time negligible relative to the total training time.
@@ 8 Discussion
Parallelism optimization. Many studies [47, 48, 49] refine parallelism strategies for deep learning models, but they fall short for multimodal LLMs due to their expansive search spaces and large cluster requirements. These methods separately optimize different strategies, resulting in inefficient parallelism Additionally, these methods generally assume data homogeneity, standardizing computations across all training samples. This assumption does not hold for multimodal LLMs due to the data heterogeneity in modality inputs of each sample. In contrast, MMScale leverages the specific training pattern of multimodal LLMs, creating a customized resource allocation problem that integrates tensor, pipeline, and data parallelism simultaneously. Besides, MMScale leverages data-aware reordering to harness the data heterogeneity.
Sequence and expert parallelism. Sequence parallelism (SP) [50] is designed to partition the training sequence into multiple subsequences for parallel training. It addresses the challenges of processing long sequences in LLMs. Expert parallelism (EP) [51], specifically devised for mixture-of-experts (MoE) LLMs [52], enables parallel training of multiple feedforward network (FFN) experts. These parallelism strategies are orthogonal to multimodal LLM training. In MMScale, both SP and EP are integrated into the training framework. MMScale treats SP and EP sizes as predefined parameters in the adaptive resource allocation optimization problem.
Disaggregating multimodal modules from LLM. In MMScale, we address model heterogeneity by treating multimodal modules and LLM as separate stages in pipeline parallelism (PP). Another approach might integrate some LLM layers with multimodal modules into the same PP stage, while distributing the remaining LLM layers to maintain balance. However, this often becomes suboptimal since it requires multimodal modules to adopt the same TP strategy as the LLM layers within the same PP stage. Such alignment degrades performance when the optimal TP strategies for multimodal modules differ from those of LLM layers. Therefore, we choose to disaggregate multimodal modules from LLM as separate PP stages to achieve better performance.
@@ 9 Related Work
LLM training. Many efforts have been made to optimize the training of LLMs from system perspectives. For LLM pretrain, Megatron-LM [17] and DeepSpeed-Megatron [53] propose customized 3D-parallelism and are de facto standards for training large LLMs. DeepSpeed-ZeRO [32] and Pytorch-FSDP [54] reduce redundant memory consumption in data parallelism. HPN [55] proposes a new dual-ToR network architecture trailed for LLM training to reduce ECMP conflict. A set of works [56, 29, 42, 43, 44] overlap the communication and computation operators in LLM training. Fault tolerance through replication and checkpoint is advanced in large training clusters by studies [29, 57]. Efforts like [58, 59, 60] further optimize recovery process in cloud spot instance scenarios. These system optimizations of LLM training are orthogonal to MMScale. They overlook the model and data heterogeneity in multimodal LLMs. MMScale also integrates several of these optimizations in training the LLM backbone.
Multimodal model training. Small multimodal models (e.g., CLIP [61] and LiT [62]) have been widely studied in recent years. Many system optimizations have been proposed to train such multimodal models efficiently. DistMM [63] tackles model heterogeneity by introducing modality-aware placement and partitioning to evenly distribute workload. GraphPipe [64] presents graph pipeline parallelism, addressing graph dependencies in multimodal models to minimize pipeline bubbles. Yet, these advancements primarily enhance small multimodal models trained on tens of GPUs. They fall short for scaling up to meet the demands of integrating multimodal models with LLMs which necessitate training across thousands of GPUs. This gap underpins the motivation behind MMScale, designed to meet the unique challenges of multimodal large language model training.
Multimodal LLM serving. LLM serving has been widely studied in recent years. Orca [65], FastServe [66], and VTC [67] propose iteration-level scheduling algorithms to improve the serving quality. DistServe [68] and Splitwise [69] propose disaggregated serving for prefill and decoding to improve the serving throughput. vLLM [70], RAGCache [71], and SGLang [72] propose prefix caching to reuse the KV tensors. However, the serving of multimodal LLMs remains under-explored. MMScale’s core insights (e.g., disaggregated modules) are applicable to multimodal LLMs serving. Our future work will delve into the specific challenges posed by multimodal LLMs serving systems.
@@ 10 Conclusion
We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs. We identify the key challenges in training multimodal LLMs, i.e., model heterogeneity and data heterogeneity. MMScale introduces novel adaptive resource allocation to eliminate model heterogeneity and data-aware reordering to mitigate data heterogeneity. We evaluate MMScale on production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on throughput.

### v2

@@ 8.3 Case Study
In this subsection, we first evaluate DistTrain under different frozen training settings. We then evaluate the overhead of CPU preprocessing disaggregation and the effectiveness of StepCCL in mitigating the TP overhead.
@@ 8.3.1 Frozen Training
We conduct a frozen training experiment under four specific training settings: complete module freezing, exclusive encoder training, exclusive LLM training, and exclusive generator training. We keep training the projectors. In these scenarios, frozen modules neither compute weight gradients nor update weights. All other experimental setup aligns with those detailed in §8.2. We evaluate the MFU and training throughput (TPT.) of DistTrain compared to Megatron-LM. The experimental results are presented in Figure 19 and Figure 20. DistTrain consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4-2.9 higher MFU and 1.2-2.9 higher training throughput. This pronounced performance gap underscores the challenges posed by Megatron-LM’s rigid model orchestration in complex training environments. In contrast, DistTrain adaptively adjusts model orchestration based on training settings and consistently achieves high resource utilization.
@@
Overhead of data preprocessing. We conduct an experiment to evaluate the overhead of data preprocessing, including decompression and reordering. Setting the DP size to one, we measure the average data preprocessing time per iteration on the GPU training side. We then compare data preprocessing time with and without CPU preprocessing disaggregation and use varying numbers of images and different image resolutions for one training iteration. The results, depicted in Figure 21, indicate that the disaggregation significantly reduces preprocessing time from seconds to milliseconds. The first parameter in the x-axis represents the number of images, while the second parameter denotes the image resolution. In production training (§8.1), iteration times range from seconds to tens of seconds. Preprocessing overhead, initially counted in seconds, significantly interferes with training. With disaggregated data preprocessing, the overhead reduces to milliseconds, which is negligible relative to total iteration time.
@@ 9 Discussion
Parallelism optimization. Many studies [47, 48, 49] refine parallelism strategies for deep learning models, but they fall short for multimodal LLMs due to their expansive search spaces and large cluster requirements. These methods separately optimize different strategies, resulting in inefficient parallelism Additionally, these methods generally assume data homogeneity, standardizing computations across all training samples. This assumption does not hold for multimodal LLMs due to the data heterogeneity in modality inputs of each sample. In contrast, DistTrain leverages the specific training pattern of multimodal LLMs, creating a customized model orchestration problem that integrates tensor, pipeline, and data parallelism simultaneously. Besides, DistTrain leverages disaggregated data reordering to harness the data heterogeneity.
Sequence and expert parallelism. Sequence parallelism (SP) [50] is designed to partition the training sequence into multiple subsequences for parallel training. It addresses the challenges of processing long sequences in LLMs. Expert parallelism (EP) [51], specifically devised for mixture-of-experts (MoE) LLMs [52], enables parallel training of multiple feedforward network (FFN) experts. These parallelism strategies are orthogonal to multimodal LLM training. In DistTrain, both SP and EP are integrated into the training framework. DistTrain treats SP and EP sizes as predefined parameters in the disaggregated model orchestration optimization problem.
@@ 10 Related Work
LLM training. Many efforts have been made to optimize the training of LLMs from system perspectives. For LLM pretrain, Megatron-LM [17] and DeepSpeed-Megatron [53] propose customized 3D-parallelism and are de facto standards for training large LLMs. DeepSpeed-ZeRO [32] and Pytorch-FSDP [54] reduce redundant memory consumption in data parallelism. HPN [55] proposes a new dual-ToR network architecture trailed for LLM training to reduce ECMP conflict. A set of works [56, 29, 42, 43, 44] overlap the communication and computation operators in LLM training. Fault tolerance through replication and checkpoint is advanced in large training clusters by studies [29, 57]. Efforts like [58, 59, 60] further optimize recovery process in cloud spot instance scenarios. These system optimizations of LLM training are orthogonal to DistTrain. They overlook the model and data heterogeneity in multimodal LLMs. DistTrain also integrates several of these optimizations in training the LLM backbone.
Multimodal model training. Small multimodal models (e.g., CLIP [61] and LiT [62]) have been widely studied in recent years. Many system optimizations have been proposed to train such multimodal models efficiently. DistMM [63] tackles model heterogeneity by introducing modality-aware placement and partitioning to evenly distribute workload. GraphPipe [64] presents graph pipeline parallelism, addressing graph dependencies in multimodal models to minimize pipeline bubbles. Yet, these advancements primarily enhance small multimodal models trained on tens of GPUs. They fall short for scaling up to meet the demands of integrating multimodal models with LLMs which necessitate training across thousands of GPUs. This gap underpins the motivation behind DistTrain, designed to meet the unique challenges of multimodal large language model training.
Multimodal LLM serving. LLM serving has been widely studied in recent years. Orca [65], FastServe [66], and VTC [67] propose iteration-level scheduling algorithms to improve the serving quality. DistServe [68] and Splitwise [69] propose disaggregated serving for prefill and decoding to improve the serving throughput. vLLM [70], RAGCache [71], and SGLang [72] propose prefix caching to reuse the KV tensors. However, the serving of multimodal LLMs remains under-explored. DistTrain’s core insights (e.g., disaggregation) are applicable to multimodal LLMs serving. Our future work will delve into the specific challenges posed by multimodal LLMs serving systems.
@@ 11 Conclusion
We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs. We identify the key challenges in training multimodal LLMs, i.e., model heterogeneity and data heterogeneity. DistTrain introduces disaggregated training to address these challenges, including disaggregated model orchestration to address model heterogeneity and disaggregated data reordering to address data heterogeneity. We evaluate DistTrain on production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on throughput.